TutoriaIs:

๐š‚๐š˜๐š•๐šŸ๐š’๐š—๐š ๐š๐šŠ๐š๐š’๐šŒ๐šŠ๐š• ๐™ด๐šš๐šž๐šŠ๐š๐š’๐š˜๐š—๐šœ; ๐’ฏ๐“Š๐“‰๐‘œ๐“‡๐’พ๐’ถ๐“: ~๐˜Œ๐˜น๐˜ฑ๐˜ญ๐˜ข๐˜ช๐˜ฏ๐˜ฆ๐˜ฅ~

TutoriaIs:

๐˜Œ๐˜น๐˜ข๐˜ฎ๐˜ฑ๐˜ญ๐˜ฆ ๐˜Œ๐˜ฒ๐˜ถ๐˜ข๐˜ต๐˜ช๐˜ฐ๐˜ฏ: (๐Œ๐จ๐ซ๐ž ๐‚๐ก๐š๐ฅ๐ฅ๐ž๐ง๐ ๐ข๐ง๐ ) \[(\sqrt{3x-8}+1)^2=(\sqrt{x+5})^2\]

TutoriaIs:

๐’๐ข๐ง๐œ๐ž ๐ญ๐ก๐ž ๐ซ๐ข๐ ๐ก๐ญ ๐ฌ๐ข๐๐ž ๐ข๐ฌ ๐ฌ๐ช๐ฎ๐š๐ซ๐ž๐, ๐ฒ๐จ๐ฎ ๐œ๐š๐ง ๐ ๐จ ๐š๐ก๐ž๐š๐ ๐š๐ง๐ ๐ญ๐š๐ค๐ž ๐จ๐ฎ๐ญ ๐ญ๐ก๐ž ๐ฌ๐ช๐ฎ๐š๐ซ๐ž ๐ซ๐จ๐จ๐ญ.

TutoriaIs:

๐๐ฒ ๐๐จ๐ข๐ง๐  ๐ญ๐ก๐š๐ญ, ๐ฒ๐จ๐ฎ ๐œ๐š๐ง ๐ญ๐ก๐ž๐ง ๐ฌ๐ž๐ž ๐ฐ๐ก๐š๐ญ ๐ญ๐ก๐ž ๐ž๐ช๐ฎ๐š๐ญ๐ข๐จ๐ง ๐ฅ๐จ๐จ๐ค๐ฌ ๐ฅ๐ข๐ค๐ž ๐ง๐จ๐ฐ. \[(\sqrt{3x-8}+1)^2=x+5\]

TutoriaIs:

๐’๐ข๐ง๐œ๐ž ๐ญ๐ก๐ž ๐ฅ๐ž๐Ÿ๐ญ ๐ฌ๐ข๐๐ž ๐ข๐ฌ ๐ฌ๐ช๐ฎ๐š๐ซ๐ž๐, ๐ญ๐ฎ๐ซ๐ง ๐ญ๐ก๐š๐ญ ๐ข๐ง๐ญ๐จ ๐š ๐›๐ข๐ง๐จ๐ฆ๐ข๐š๐ฅ: \[(\sqrt{3x-8}+1)~(\sqrt{3x-8}+1)=x+5\]

TutoriaIs:

๐๐จ๐ฐ, ๐ฅ๐ž๐ญโœ๐ฌ ๐ญ๐ซ๐ฒ ๐ญ๐จ ๐ ๐ž๐ญ ๐ญ๐ก๐ž ๐ฌ๐ช๐ฎ๐š๐ซ๐ž ๐ซ๐จ๐จ๐ญ ๐›๐ฒ-๐ข๐ญ๐ฌ๐ž๐ฅ๐Ÿ. \[3x-7+2\sqrt{3x-8}=x+5\] ๐’๐ฎ๐›๐ญ๐ซ๐š๐œ๐ญ ๐Ÿ‘๐ฑ ๐š๐ง๐ ๐š๐๐ ๐Ÿ• ๐ญ๐จ ๐ž๐š๐œ๐ก ๐ฌ๐ข๐๐ž. ๐๐จ๐ฐ, ๐ฐ๐žโœ๐ซ๐ž ๐ ๐จ๐ข๐ง๐  ๐ญ๐จ ๐›๐š๐ฅ๐š๐ง๐œ๐ž ๐ญ๐ก๐ž ๐ž๐ช๐ฎ๐š๐ญ๐ข๐จ๐ง ๐›๐ฒ ๐๐ข๐ฏ๐ข๐๐ข๐ง๐  ๐ž๐ฏ๐ž๐ซ๐ฒ๐ญ๐ก๐ข๐ง๐  ๐›๐ฒ ๐Ÿ. \[\frac{ 2\sqrt{3x-8} =-2x+12}{ 2 }\] ๐–๐ก๐ข๐œ๐ก ๐ญ๐ก๐ž๐ง ๐ž๐ช๐ฎ๐š๐ฅ๐ฌ: \[\sqrt{3x-8}=-x+6\] ๐๐จ๐ฐ, ๐ฐ๐žโœ๐ซ๐ž ๐ ๐จ๐ข๐ง๐  ๐ญ๐จ ๐ฌ๐ช๐ฎ๐š๐ซ๐ž ๐ž๐ฏ๐ž๐ซ๐ฒ๐ญ๐ก๐ข๐ง๐  ๐ญ๐จ ๐ ๐ž๐ญ ๐ซ๐ข๐ ๐จ๐Ÿ ๐ญ๐ก๐ž ๐ฌ๐ช๐ฎ๐š๐ซ๐ž ๐ซ๐จ๐จ๐ญ ๐จ๐ง ๐ญ๐ก๐ž ๐ฅ๐ž๐Ÿ๐ญ ๐ฌ๐ข๐๐ž. \[(\sqrt{3x-8})^2=(-x+6)^2\] ๐’๐จ, ๐ž๐š๐œ๐ก ๐ฌ๐ข๐๐ž ๐ฐ๐ข๐ฅ๐ฅ ๐ง๐จ๐ญ ๐›๐ž ๐ฌ๐ช๐ฎ๐š๐ซ๐ž๐ ๐š๐ง๐ฒ๐ฆ๐จ๐ซ๐ž. ๐๐จ๐ฐ, ๐ข๐ญโœ๐ฌ ๐ž๐š๐ฌ๐ข๐ž๐ซ ๐ญ๐จ ๐ฌ๐ž๐ญ ๐ญ๐ก๐ž๐ฆ ๐ž๐ช๐ฎ๐š๐ฅ ๐ญ๐จ ๐ž๐š๐œ๐ก ๐จ๐ญ๐ก๐ž๐ซ. \[3x-8=(-x+6)^2\] ๐“๐ฎ๐ซ๐ง ๐ญ๐ก๐ž ๐ž๐ช๐ฎ๐š๐ญ๐ข๐จ๐ง ๐จ๐ง ๐ญ๐ก๐ž ๐ซ๐ข๐ ๐ก๐ญ ๐ข๐ง๐ญ๐จ ๐š ๐›๐ข๐ง๐จ๐ฆ๐ข๐š๐ฅ: \[3x-8(-x+6)(-x+6)\] ๐Œ๐ฎ๐ฅ๐ญ๐ข๐ฉ๐ฅ๐ฒ ๐ญ๐ก๐ž ๐ฉ๐š๐ซ๐ž๐ง๐ญ๐ก๐ž๐ฌ๐ž๐ฌ: \[3x-8=x^2-6x-6x+36\] ๐’๐ฎ๐›๐ญ๐ซ๐š๐œ๐ญ ๐Ÿ‘๐ฑ ๐š๐ง๐ ๐š๐๐ ๐Ÿ– ๐ญ๐จ ๐ž๐š๐œ๐ก ๐ฌ๐ข๐๐ž ๐จ๐Ÿ ๐ญ๐ก๐ž ๐ž๐ช๐ฎ๐š๐ญ๐ข๐จ๐ง: \[0=x^2-15x+44\] ๐๐จ๐ฐ, ๐ฐ๐ž ๐œ๐š๐ง ๐Ÿ๐š๐œ๐ญ๐จ๐ซ ๐ญ๐ก๐ž ๐Ÿ’๐Ÿ’. \[0=(x-11)(x-4)\] ๐๐จ๐ฐ, ๐ฐ๐žโœ๐ฏ๐ž ๐Ÿ๐จ๐ฎ๐ง๐ ๐จ๐ฎ๐ซ ๐š๐ง๐ฌ๐ฐ๐ž๐ซ. :) ๐๐ฎ๐ญ, ๐ฅ๐ž๐ญโœ๐ฌ ๐ฆ๐š๐ค๐ž ๐ฌ๐ฎ๐ซ๐ž ๐ฐ๐žโœ๐ฏ๐ž ๐ ๐จ๐ญ ๐ญ๐ก๐ž ๐œ๐จ๐ซ๐ซ๐ž๐œ๐ญ ๐š๐ง๐ฌ๐ฐ๐ž๐ซ ๐›๐ž๐œ๐š๐ฎ๐ฌ๐ž ๐ฌ๐จ๐ฆ๐ž๐ญ๐ข๐ฆ๐ž๐ฌ, ๐ฒ๐จ๐ฎ ๐œ๐จ๐ฎ๐ฅ๐โœ๐ฏ๐ž ๐๐จ๐ง๐ž ๐ฆ๐š๐ญ๐ก ๐ข๐ง๐œ๐จ๐ซ๐ซ๐ž๐œ๐ญ๐ฅ๐ฒ. ๐’๐จ, ๐ฅ๐ž๐ญโœ๐ฌ ๐ข๐ง๐ฌ๐ญ๐š๐ฅ๐ฅ ๐ญ๐ก๐ž๐ฆ ๐ข๐ง๐ญ๐จ ๐ญ๐ก๐ž ๐ž๐ช๐ฎ๐š๐ญ๐ข๐จ๐ง. \[\sqrt{3x-8}+1=\sqrt{x=5}\] ๐๐ฒ ๐ข๐ง๐ฌ๐ญ๐š๐ฅ๐ฅ๐ข๐ง๐  ๐ญ๐ก๐ž๐ฆ, ๐ฒ๐จ๐ฎ ๐Ÿ๐ข๐ง๐ ๐ญ๐ก๐š๐ญ ๐Ÿ” ๐ข๐ฌ๐งโœ๐ญ ๐ž๐ช๐ฎ๐š๐ฅ ๐Ÿ’, ๐ฌ๐จ ๐Ÿ๐Ÿ ๐๐จ๐ž๐ฌ๐งโœ๐ญ ๐ฐ๐จ๐ซ๐ค. ๐’๐จ ๐ข๐ง๐ฌ๐ญ๐ž๐š๐, ๐ฅ๐ž๐ญโœ๐ฌ ๐ข๐ง๐ฌ๐ญ๐š๐ฅ๐ฅ ๐ญ๐ก๐ž ๐Ÿ’. ๐“๐ก๐ž๐ง ๐ฐ๐ž ๐Ÿ๐ข๐ง๐ ๐ญ๐ก๐š๐ญ ๐Ÿ+๐Ÿ=๐Ÿ‘, ๐ฌ๐จ ๐Ÿ’ ๐ฐ๐จ๐ซ๐ค๐ฌ. \[x=4\] ๐ˆ๐Ÿ ๐ฒ๐จ๐ฎโœ๐ ๐ฅ๐ข๐ค๐ž ๐š๐ง๐ฒ ๐ฆ๐จ๐ซ๐ž ๐ก๐ž๐ฅ๐ฉ, ๐ฃ๐ฎ๐ฌ๐ญ ๐œ๐จ๐ง๐ญ๐š๐œ๐ญ ๐ฆ๐ž.